OpenStudy (anonymous):

Faith observed 16 bugs while doing research for her science project. She observed insects and spiders, which had a total of 114 legs. Bearing in mind that insects have 6 legs and spiders have 8 legs, how many insects did she observe? Faith observed 16 bugs while doing research for her science project. She observed insects and spiders, which had a total of 114 legs. Bearing in mind that insects have 6 legs and spiders have 8 legs, how many insects did she observe? @Mathematics

OpenStudy (saifoo.khan):

Make two equations.

OpenStudy (anonymous):

I dont understand tis at all. How do you set up an equation for this?

OpenStudy (saifoo.khan):

Look, the best way to solve these is to break the question into parts.. the first sentence says, they observed 16 bugs... Let spiders be x. And insects be y. there were total 16 bugs.. so, x+y = 16

OpenStudy (anonymous):

so then the next equation would be 2x + y = 144 right?

OpenStudy (saifoo.khan):

moving on.. insects has 6 legs. while spider has 8 legs. so, 8x + 6y = 114.

OpenStudy (anonymous):

ooo ok thank you i get it now!

OpenStudy (saifoo.khan):

YAY! =D